The sum of age of father and son is 75.If the father and his son both live until the son becomes as old as his father is now,the sum of the age will be 107.
Answer by ankor@dixie-net.com(22740)

let f = father's age
let s = son's age
:
The sum of age of father and son is 75.
f + s = 75
the difference in their ages
d = f - s
:
If the father and his son both live until the son becomes as old as his father is now, the sum of the age will be 107.
:
(f + d) + (s + d) = 107
f + s + 2d = 107
we know that f + s = 75, from the 1st statement
75 + 2d = 107
2d = 107 - 75
2d = 32
d = 16 yrs is the their age difference (a young father)
;
f = s + 16
:
f + s = 75
(s+16) + s = 75
2s = 75 - 16
2s = 59
s = 59/2
s = 29.5 yrs is the son's age
then
29.5+16 = 45.5 yrs is father's age
:
:
See if this works:
29.5 + 45.5 = 75
and
(29.5+16) + (45.5+16) =
45.5 + 61.5 = 107